OpenStudy (anonymous):

A straight segment of wire has a length of 24 cm and carries a current of 5 A. It is oriented at right angles to a magnetic field of 0.6 T. What is the magnitude of the magnetic force on this segment of wire?

OpenStudy (irishboy123):

use F = B I L, yes?

OpenStudy (anonymous):

My professor did a really bad job explaining this chapter, so I am still very unsure of what to do

OpenStudy (irishboy123):

where do you want to start? Lorentz force equation? or just what the letters stand for or what magnetism means/ does?

OpenStudy (anonymous):

What the letters stand for and then the force equation, please.

OpenStudy (michele_laino):

the requested magnitude F of the force acting on your segment is given by the subsequent formula: \[\large F = B \times I \times l\] where B is the magnetic field, I is the current, and l is the length of your segment

OpenStudy (michele_laino):

hint: \[\large F = B \times I \times l = 0.6 \times 5 \times 0.24 = ...?\]

OpenStudy (michele_laino):

since l= 24 cm = 0.24 meters

OpenStudy (anonymous):

Thanks @IrishBoy123 and @Michele_Laino

OpenStudy (michele_laino):

Thanks! @GVSU2016